Answer all three questions with showing proof of your work for full credit. Round all answers to one decimal place.

Assume that 2018 is the base year. The quantity consumed in each basket is ( 3 Cell Phones, 4 Shoes)

  • What is the CPI in 2019?
  • What is the CPI in 2020?
  • What is the inflation rate from 2018 to 2019? 

 

YearPrice of Cell PhonesPrice of Shoes2018$500$852019$575$892020$625$95


1
Expert's answer
2021-04-20T12:46:33-0400

In year 2018,

Price of cell phone=$500

Price of shoes=$85

In year 2019,

Price of cell phone=$575

Price of shoes=$89

In year 2020

Price of cell phone=$625

Price of shoes=$95

Cost of consumption basket in 2018=C1

Cost of consumption basket in year(2019) =C2

Cost of consumption basket in year(2020) =C3

"C1=(3\times500+4\times85)" =1840

"C2=(3\times575+5\times89)=2081"

"C3=(3\times625+4\times95)=2255"

Let CPI in 2018 =D1

CPI in 2019=D2

CPI in 2020=D3

"CPI=\frac{Cost of consumption in current year}{cost of consumption in base year}"

CPI in 2018=1

CPI in 2019=2081/1840=1.13

CPI in 2020=2255/1840=1.22

Inflation rate is 22%